A right triangle is a triangle with one angle measuring 90 degree. In a right triangle, the sides are related by the Pythagorean Theorem, c^2 = a^2 + b^2 where c is the hypotenuse (the side opposite the 90 degree angle). Find the hypotenuse when the other 2 sides' measurements are 3 feet and 4 feet.
Answer by checkley71(8403)

C^2=3^2+4^2
C^2=9+16
C^2=25
C=SQRT25
C=5 IS THE HYPOTENUSE